# Computing Chebyshev polynomials using the complex Remez algorithm

### v2

@@ 4 Discussion
4.1 Widom factors
The Widom factors computed in Section 3 are computed to a high degree of accuracy. We believe that Tang’s algorithm can be very useful in getting suggested behavior regarding the Widom factors corresponding to a set. This method has previously been applied in [8] where a result on the limits of Widom factors – first conjectured using numerical experiments – was resolved theoretically. The conjecture whose validity we wish to argue for is the following.
Conjecture 1.
Let $\mathsf{E}\subset\mathbb{C}$ denote the closure of a Jordan domain with piecewise analytic boundary where none of the singularities of $\partial\mathsf{E}$ are cusp points. Then
$\lim_{n\rightarrow\infty}\mathcal{W}_{n}(\mathsf{E})=1.$
4.1.1 Regular polygon
We begin by discussing the Widom factors computed for the regular polygons. As we previously remarked, it is known that these are bounded by $2$ due to the inherent convexity of the set but apart from this bound, not much is known. The plots in Figures 1(a)-1(d) clearly suggests that $\{\mathcal{W}_{n}(\mathsf{E})\}$ is monotonically decreasing in $n$ if $n>2$ and $\mathsf{E}$ is an $m$-gon. Furthermore, it seems to be the case that the Widom factors converge to $1$. This is in accordance with Conjecture 1. The computations clearly suggests that there is differing behavior between Chebyshev polynomials and Faber polynomials corresponding to the regular $m$-gon in terms of their supremum value. Indeed, by [31, Theorem II.2.1], we see that if $\mathsf{E}_{m}$ is an $m$-gon with corners at $\exp(\frac{2\pi ik}{m})$ then
@@ 4.1.1 Regular polygon
as $n\rightarrow\infty$ for $k=0,1,2\dotsc,m-1$. In conclusion, we see that
@@
We remark that the dotted lines visible in Figures 1(a)-1(d) represent the value $(2+m)/m$. If we choose to believe that $\mathcal{W}_{n}(\mathsf{E}_{m})$ decrease monotonically for $n\geq 2$ then as Figures 1(a)-1(d) illustrate, the norms of the Chebyshev polynomials are significantly smaller.
Based on these considerations, the Faber polynomials corresponding to the regular polygons presumably do not provide good enough estimates as trial polynomials to determine the limits of the Widom factors. In short, we believe that the sequence $\{\mathcal{W}_{n}(\mathsf{E}_{m})\}$ decreases monotonically if $n\geq 2$ and that the limit is $1$ as $n\rightarrow\infty$.
One approach in proving that the limit value is $1$ is to analyze some well-suited family of trial polynomials whose normalized norms converge to $1$. How to construct such a family is not immediately clear to us. Under the assumption that $\lim_{n\rightarrow\infty}\mathcal{W}_{n}(\mathsf{E}_{m})=1$ holds this would not constitute the only example where the Faber polynomials are ill-suited trial polynomials for determining the detailed behavior of $\mathcal{W}_{n}(\mathsf{E})$. In the extreme case, an example of Clunie [13] further studied by Suetin [43, p. 179] and Gaier [17] illustrates the existence of a quasi-disk $\mathsf{E}$ such that the quantity
@@
is unbounded in $n$ along some sparse subsequence. In comparison [4, Theorem 1] shows that $\mathcal{W}_{n}(\mathsf{E})$ is still bounded in this case.
4.1.2 Hypocycloid
Recall that $\mathsf{H}_{m}$ denotes the $m$-cusped hypocycloid defined in (19). Since $\mathsf{H}_{m}$ is piecewise analytic away from the cusp points which are outward pointing, [49, Theorem 2.1] can be applied to deduce that $\mathcal{W}_{n}(\mathsf{H}_{m})$ is bounded. The Faber polynomials again seem ill-suited in order draw conclusions concerning the precise behavior of the Widom factors in this case since it is shown in [24] that
@@ 4.1.2 Hypocycloid
holds or not. The numerical experiments illustrated in Figures 2(a)-2(d) paint a richer picture. Again, it seems likely that the sequence $\mathcal{W}_{n}(\mathsf{H}_{m})$ decreases monotonically, suggesting that the sequence $\mathcal{W}_{n}(\mathsf{H}_{m})$ has a limit as $n\rightarrow\infty$. In comparison to the Widom factors of the regular $m$-gons, the decay appears to be slower in this case. We find it reasonable to assume that that $\lim_{n\rightarrow\infty}\mathcal{W}_{n}(\mathsf{H}_{m})$ – if it exists – should be smaller than $2$ due to the monotonicity pattern and the values computed in Table 2. We find it difficult to say whether the correct conjecture is that the sequence converges to the theoretical minimal value of $1$ since the decay seems to be slow. For this reason we believe that “outward pointing cusps” should be excluded from Conjecture 1 since it is not clear even in the case of the $m$-cusped hypocycloid if the associated Widom factors asymptotically saturates (3).
4.1.3 Circular lune
Recall that $\mathsf{C}_{\alpha}$, defined in (20), denotes the circular lune with vertices at ±$\pm\alpha$ and exterior angle $\pi\alpha$. Based on the plots in Figures 3(a) and 3(b) together with the computations in Table 3, it seems likely that the Widom factors corresponding to $\mathsf{C}_{\alpha}$ converges to $1$. It is interesting to note that when the set is convex then the whole sequence $\mathcal{W}_{n}(\mathsf{C}_{\alpha})$ appears to be monotonically decreasing, see Figure 3(b). On the other hand, if $\alpha\in(0,1)$ then two distinct monotonically decreasing subsequences of $\mathcal{W}_{n}(\mathsf{C}_{\alpha})$ emerge based on the parity of the degrees. We believe that the sequence $\{\mathcal{W}_{2n+l}(\mathsf{C}_{\alpha})\}_{n}$ is monotonically decreasing to $1$ for fixed $l\in\{0,1\}$ if $\alpha\in(0,2)$. The case that $\alpha=2$ is excluded for it is classical that $\mathcal{W}_{n}(\mathsf{C}_{2})=2$ for any value of $n$. Also classical is the fact that $\mathcal{W}_{n}(\mathsf{C}_{1})=1$. We believe that this example shows that for a nice enough bounding curve, it is not necessary that the set is convex for the sequence of Widom factors to converge to the theoretical minimal value. This also motivates our quite general formulation of Conjecture 1.
4.2 Motivating the Faber connection
The Chebyshev polynomials and Faber polynomials will both exhibit similar symmetric structure as the corresponding underlying set. To see this, one should compare Lemma 1 to [21, Theorem 2.2] or [22, Theorem 2.1]. This comparison is essentially encapsulated in the following simple lemma.
Lemma 2.
If $\mathsf{E}$ is invariant under rotations of $2\pi/m$ then both the Chebyshev polynomial and Faber polynomial of respective degrees $nm+l$ are polynomials in $z^{m}$ multiplied by the factor $z^{l}$ for $l\in\{0,1,\dotsc,m-1\}$. In particular,
@@ 4.2 Motivating the Faber connection
We already saw that $T_{l}^{\mathsf{E}}=z^{l}$ in this case in Lemma 1. The result follows from the fact that if $\Phi$ is the conformal map from (6) then
@@
if $l\in\{0,1,\dotsc,m-1\}$. In the special case where $n=0$ we see that $F_{l}^{\mathsf{E}}(z)=z^{l}$. ∎
Of course Lemma 2 has more to do with the rotational symmetry of a set than any other property. It does, however, give several easy examples where the two families of polynomials overlap.
@@
for $z\in\mathsf{E}^{r}:=\{\zeta:|\Phi(\zeta)|=r\}$. If $[\operatorname{Cap}(\mathsf{E})\Phi(z)]^{n}$ was a polynomial of degree $n$ then it would follow from (3) that it would coincide with the corresponding Chebyshev polynomial. We have already seen examples of this when studying lemniscates. Although this is rarely the case, we observe that $F_{n}^{\mathsf{E}}$ will be an increasingly good candidate for obtaining relatively small maximum values on $\mathsf{E}^{r}$ as $r\rightarrow\infty$. For a fixed degree $n$, $F_{n}^{\mathsf{E}}$ will be asymptotically minimal on $\mathsf{E}^{r}$ in the sense that
@@
$T_{n}^{\mathsf{E}^{r}}\rightarrow F_{n}^{\mathsf{E}}$
as $r\rightarrow\infty$ to hold in general. Based on the numerical data illustrated in Figure 4 this is clearly hinted upon for these specific domains. We therefore make the following conjecture.
Conjecture 2.
Let $\mathsf{E}$ denote a connected compact set with simply connected complement and let $\Phi:\mathbb{C}\setminus\mathsf{E}\rightarrow\{z:|z|>1\}$ denote the conformal map of the form
$\Phi(z)=\operatorname{Cap}(\mathsf{E})^{-1}z+a_{0}+a_{-1}z^{-1}+\cdots.$
If $\mathsf{E}^{r}=\{z:|\Phi(z)|=r\}$ then
$\lim_{r\rightarrow\infty}T_{n}^{\mathsf{E}^{r}}=F_{n}^{\mathsf{E}}.$
We find the data presented in Figure 4 convincing in suggesting the validity of Conjecture 2 for these specific types of sets and remark that similar patterns have materialized for any other combination of degrees and sets that we have considered. In the general case it is clear that $\mathsf{E}^{r}$ will be an analytic curve for $r>1$ and hence the regularity of the boundary of $\mathsf{E}$ is perhaps of less importance since the Faber polynomials corresponding to $\mathsf{E}$ are the same as the ones corresponding to $\mathsf{E}^{r}$. We stress again the fact that the algorithm outputs polynomials $P_{n}$ such that $\|P_{n}\|_{\mathsf{E}}-\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}$ is small. This is not exactly the same as saying that $\|P_{n}-T_{n}^{\mathsf{E}}\|_{\infty}$ is small with $\|\cdot\|_{\infty}$ defined in (24). What is true, is that for a fixed $n$, $\|P_{n}\|_{\mathsf{E}}\rightarrow\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}$ implies that $\|P_{n}-T_{n}^{\mathsf{E}}\|_{\infty}\rightarrow 0$. The computations remain consistent throughout. No matter how close we approximate the minimal norm, the behavior as suggested in Figure 4 remains.
4.3 Zero distributions
@@
It is shown in [40] that the zeros corresponding to the closure of a Jordan domain stay away from the boundary precisely when the bounding curve is analytic. As such we see that in all of our examples, except for the cases of lemniscates $\{z:|P(z)|=r\}$ with analytic boundary, the zeros should approach some part of the boundary. From [11, Theorem 1.1] we gather that every “corner point” on the respective sets $\mathsf{C}_{\alpha}$, $\mathsf{E}_{m}$ and $\mathsf{H}_{m}$ should attract zeros. This also appears to be the case, albeit, slowly for $\mathsf{C}_{1/2}$.
Predicting the behavior of zeros of extremal polynomials based on plots has proven hazardous in the past. In particular, we refer to the reader to [38] where five conjectures concerning limiting zero distributions are made very plausible using numerical plots only to be proven to be wrong using theoretical results. However, if one chooses to believe that Conjecture 2 is true then this alludes to the possibility that potential weak-star limits of $\nu(T_{n}^{\mathsf{E}})$ and $\nu(F_{n}^{\mathsf{E}})$ are related. In the cases where we boldly propose conjectures regarding weak-star convergence of $\nu(T_{n}^{\mathsf{E}})$ we emphasize that the corresponding weak-star limits of the counting measures $\nu(F_{n}^{\mathsf{E}})$ are known to have this very behavior.
4.3.1 Regular polygons
We adopt the notation $\mathsf{E}_{m}$ to denote the regular polygon with $m$ sides. As is suggested by Figures 5(a)-5(d), the zeros of $T_{n}^{\mathsf{E}_{m}}$ for low degrees appear to lie on the diagonal lines between the vertices and the origin. However, by increasing the degree it seems clear that the zeros approach the boundary. In [38] the case of Faber polynomials on $\mathsf{E}_{3}$ are discussed. Here the authors specify that for small degrees the zeros of $F_{n}^{\mathsf{E}_{3}}$ appear to distribute along the diagonals however they also note that as a consequence of [29, Theorem 1.5] at least a subsequence of $\nu(F_{n}^{\mathsf{E}_{3}})$ converges in the weak-star sense to $\mu_{\mathsf{E}_{3}}$ which is supported on the boundary. The zeros of certain $F_{n}^{\mathsf{E}_{m}}$ are illustrated in [22] and appear to behave very similar to the ones for $T_{n}^{\mathsf{E}_{m}}$ computed here. We therefore believe that the zeros should approach the boundary in the sense that (11) should hold for every compact set in the interior. This would of course also imply that
@@ 4.3.1 Regular polygons
4.3.2 Circular lune
Recall the definition of $\mathsf{C}_{\alpha}$ from (20). Based on the plot in Figure 5(e) it appears as most of the zeros approach the boundary in the case when $\alpha=1/2$ and that (11) should hold for any compact set contained in the interior. This is in fact a known result and follows from [39, Theorem 2.1]. Indeed, from there we gather that $\nu(T_{n}^{\mathsf{C}_{\alpha}})\xrightarrow{\ast}\mu_{\mathsf{C}_{\alpha}}$ as $n\rightarrow\infty$ for any $\alpha\in(0,1)$. In this sense the computed polynomials serves to confirm the predicted behaviour from theoretical results. For any value of $\alpha\in(0,1)\cup(1,2)$ it follows from [29, Theorem 1.5] that $\nu(F_{n}^{\mathsf{C}_{\alpha}})\xrightarrow{\ast}\mu_{\mathsf{C}_{\alpha}}$ along some subsequence. Again, motivated by the belief that the conjectured similarities between Chebyshev polynomials and Faber polynomials persists for $\mathsf{C}_{\alpha}$ together with the strong resemblance between the plots of zeros for Faber polynomials in [22] with the corresponding zeros of $T_{n}^{\mathsf{C}_{\alpha}}$ computed here, we suspect that

### v3

@@ 4. Discussion
4.1. Widom factors
The Widom factors computed in Section 3 are computed with an accuracy of $10^{-10}$ within Tang’s algorithm. We believe that Tang’s algorithm can be very useful in getting suggesting the behavior of the Widom factors corresponding to a set. This method has previously been applied in [9] where a result on the limits of Widom factors – first conjectured using numerical experiments – was resolved theoretically. We wish to argue for the validity that
$\lim_{n\rightarrow\infty}\mathcal{W}_{n}(\mathsf{E})=1$
holds for closures of Jordan domains with piecewise-analytic boundary.
4.1.1. Regular polygons
We begin by discussing the Widom factors computed for the regular polygons. As we previously remarked, it is known that the Widom factors are bounded by $2$. Apart from this bound, not much is known. The plots in Figures 1(a)-1(d) clearly suggests that $\{\mathcal{W}_{n}(\mathsf{E})\}$ is monotonically decreasing in $n$ if $n>2$ and $\mathsf{E}$ is an $m$-gon. The numerics suggests that the Widom factors converge to $1$ at a rate proportional to $1/n$. This observation points to a difference in the supremum norm between the Chebyshev polynomials and the Faber polynomials associated with the regular $m$-gon. Indeed, by [35, Theorem II.2.1], we see that if $\mathsf{E}_{m}$ is an $m$-gon with corners at $\exp(\frac{2\pi ik}{m})$ then
@@ 4.1.1. Regular polygons
as $n\rightarrow\infty$ for $k=0,1,2\dotsc,m-1$. In particular,
@@
We remark that the dotted lines visible in Figures 1(a)-1(d) represent the value $(2+m)/m-1$. If we choose to believe that the sequence $\{\mathcal{W}_{n}(\mathsf{E}_{m})\}$ eventually decreases monotonically for $n$ large enough then as Figures 1(a)-1(d) illustrate, the norms of the Chebyshev polynomials are significantly smaller.
Based on these considerations, the Faber polynomials corresponding to the regular polygons presumably do not provide good enough estimates as trial polynomials to determine the limits of the Widom factors. In short, we believe that the sequence $\{\mathcal{W}_{n}(\mathsf{E}_{m})\}$ eventually decreases monotonically and that the limit is $1$ as $n\rightarrow\infty$.
One approach in proving that the limit value is $1$ is to analyze some well-suited family of trial polynomials whose normalized norms converge to $1$. How to construct such a family is not immediately clear to us. Under the assumption that $\lim_{n\rightarrow\infty}\mathcal{W}_{n}(\mathsf{E}_{m})=1$ holds this would not constitute the only example where the Faber polynomials are ill-suited trial polynomials for determining the detailed behavior of $\mathcal{W}_{n}(\mathsf{E})$. In the extreme case, an example of Clunie [14] further studied by Suetin [48, p. 179] and Gaier [19] shows the existence of a quasi-disk $\mathsf{E}$ such that the quantity
@@
is unbounded in $n$ along some sparse subsequence. In comparison, [5, Theorem 1] shows that $\mathcal{W}_{n}(\mathsf{E})$ is still bounded in this case.
4.1.2. Hypocycloids
Recall that $\mathsf{H}_{m}$ denotes the $m$-cusped hypocycloid defined in (22). Since $\mathsf{H}_{m}$ is piecewise-analytic away from the cusp points which are outward pointing, [55, Theorem 2.1] can be applied to deduce that $\mathcal{W}_{n}(\mathsf{H}_{m})$ is bounded. The Faber polynomials again seem ill-suited in order draw conclusions concerning the precise behavior of the Widom factors in this case, since it is shown in [26] that
@@ 4.1.2. Hypocycloids
holds or not. The numerical experiments illustrated in Figures 2(a)-2(d) paint a richer picture. Again, it seems that the sequence $\mathcal{W}_{n}(\mathsf{H}_{m})$ converges to $1$ with a convergence rate proportional to $n^{-1/2}$ and that the sequence $\mathcal{W}_{n}(\mathsf{H}_{m})$ is monotone decreasing if $n$ is large enough. In comparison to the Widom factors of the regular $m$-gons, the decay appears to be slower in the case of hypocycloids.
4.1.3. Circular lunes
Recall that $\mathsf{C}_{\alpha}$, defined in (23), denotes the circular lune with vertices at ±$\pm\alpha$ and exterior angle $\pi\alpha$. Based on the plots in Figures 3(a) and 3(b) together with the results in Table 3, it seems likely that the Widom factors corresponding to $\mathsf{C}_{\alpha}$ converges to $1$. It is interesting to note that when the set is convex then the whole sequence $\mathcal{W}_{n}(\mathsf{C}_{\alpha})$ appears to be monotonically decreasing, see Figure 3(b). On the other hand, if $\alpha\in(0,1)$ then two distinct monotonically decreasing subsequences of $\mathcal{W}_{n}(\mathsf{C}_{\alpha})$ emerge based on the parity of the degrees. We believe that the sequence $\{\mathcal{W}_{2n+l}(\mathsf{C}_{\alpha})\}_{n}$ is monotonically decreasing to $1$ for fixed $l\in\{0,1\}$ if $\alpha\in(0,2)$. The rate of convergence appears to be different depending on the convexity properties of $\mathsf{C}_{\alpha}$. In particular, if the set is convex then the rate of decrease is proportional to $n^{-1}$ while for $\mathsf{C}_{1/2}$ the decay is slower. If $\alpha=2$ then $\mathcal{W}_{n}(\mathsf{C}_{2})=2$ for any value of $n$.
4.2. Motivating the Faber connection
The Chebyshev polynomials and Faber polynomials will both inherit symmetric structures from the underlying set. To see this, one should compare Lemma 3.1 to [23, Theorem 2.2] or [24, Theorem 2.1]. This comparison is essentially encapsulated in the following simple lemma.
Lemma 4.1.
If $\mathsf{E}$ is invariant under rotations by $2\pi/m$, $n\in\mathbb{N}$, and $l\in\{0,1,\dotsc,m-1\}$, then both $T_{nm+l}^{\mathsf{E}}$ and $F_{nm+l}^{\mathsf{E}}$ belong to
(29) $\{z^{l}Q(z^{m}):Q\text{ is a monic polynomial of degree $n$}\}.$
In particular,
@@ 4.2. Motivating the Faber connection
From Lemma 3.1 we conclude the validity of every claim made on $T_{nm+l}^{\mathsf{E}}$. The conformal map $\Phi$ from (6) necessarily satisfies
@@
if $l\in\{0,1,\dotsc,m-1\}$. In a completely analogous fashion to how Lemma 3.1 was shown, we gather that $F_{nm+l}^{\mathsf{E}}$ belongs to the set in (29). In the special case where $n=0$ we see that $F_{l}^{\mathsf{E}}(z)=z^{l}=T_{l}^{\mathsf{E}}(z)$. ∎
Of course Lemma 4.1 has more to do with the rotational symmetry of a set than any other property. It does, however, give several easy examples where the two families of polynomials overlap.
@@
for $z\in\mathsf{E}^{r}\coloneqq\{\zeta:|\Phi(\zeta)|=r\}$. If $[\operatorname{Cap}(\mathsf{E})\Phi(z)]^{n}$ was a polynomial of degree $n$ then it would follow from (3) that it would coincide with the corresponding Chebyshev polynomial. We have already seen examples of this when studying lemniscates. Although this is rarely the case, we observe that $F_{n}^{\mathsf{E}}$ will be an increasingly good candidate for obtaining relatively small maximum values on $\mathsf{E}^{r}$ as $r\rightarrow\infty$. For a fixed degree $n$, $F_{n}^{\mathsf{E}}$ will be asymptotically minimal on $\mathsf{E}^{r}$ in the sense that
@@
(30) $T_{n}^{\mathsf{E}^{r}}\rightarrow F_{n}^{\mathsf{E}}$
as $r\rightarrow\infty$ to hold in general. Based on the numerical data illustrated in Figure 4 this is clearly hinted upon for these specific domains.
We remark that similar patterns have materialized for any other combination of degrees and sets that we have considered. In the general case, it is clear that $\mathsf{E}^{r}$ will be an analytic curve for $r>1$ and hence the regularity of the boundary of $\mathsf{E}$ is perhaps of less importance since the Faber polynomials corresponding to $\mathsf{E}$ are the same as the ones corresponding to $\mathsf{E}^{r}$. We stress the fact that the algorithm outputs polynomials $P_{n}$ such that $\|P_{n}\|_{\mathsf{E}}-\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}$ is small. This is not exactly the same as saying that $\|P_{n}-T_{n}^{\mathsf{E}}\|_{\infty}$ is small with $\|\cdot\|_{\infty}$ defined in (27). What is true, is that for a fixed $n$, $\|P_{n}\|_{\mathsf{E}}\rightarrow\|T_{n}^{\mathsf{E}}\|_{\mathsf{E}}$ implies that $\|P_{n}-T_{n}^{\mathsf{E}}\|_{\infty}\rightarrow 0$. The computations remain consistent throughout. No matter how close we approximate the minimal norm, the behavior as suggested in Figure 4 remains.
4.3. Zero distribution
@@
It is shown in [45] that the zeros of Chebyshev polynomials corresponding to the closure of a Jordan domain stay away from the boundary precisely when the bounding curve is analytic. As such we see that in all of our examples, except for the cases of lemniscates $\{z:|P(z)|=r\}$ with analytic boundary, the zeros should approach some part of the boundary. From [12, Theorem 1.1] we gather that every “corner point” on the respective sets $\mathsf{C}_{\alpha}$, $\mathsf{E}_{m}$ and $\mathsf{H}_{m}$ attract zeros. This also appears to be the case, albeit, slowly for $\mathsf{C}_{1/2}$.
Predicting the behavior of zeros of extremal polynomials based on plots has proven hazardous in the past. In particular, we refer to the reader to [43] where five conjectures concerning limiting zero distributions are made very plausible using numerical plots only to be proven to be wrong using theoretical results.
4.3.1. Regular polygons
We adopt the notation $\mathsf{E}_{m}$ to denote the regular polygon with $m$ sides. As is suggested by Figures 5(a)-5(d), the zeros of $T_{n}^{\mathsf{E}_{m}}$ for low degrees appear to lie on the diagonal lines between the vertices and the origin. However, by increasing the degree it seems clear that the zeros approach the boundary in all cases but the equilateral triangle as seen in 5(a). In [43], the case of Faber polynomials on $\mathsf{E}_{3}$ are discussed. Here the authors specify that for small degrees the zeros of $F_{n}^{\mathsf{E}_{3}}$ appear to distribute along the diagonals, however, they also note that as a consequence of [32, Theorem 1.5] at least a subsequence of $\nu(F_{n}^{\mathsf{E}_{3}})$ converges in the weak-star sense to $\mu_{\mathsf{E}_{3}}$ which is supported on the boundary. We find it intriguing to determine whether the zeros of the Chebyshev polynomials on $\mathsf{E}_{3}$ remain on the diagonal for all values of $n$ as this contrasts the behavior of Faber polynomials in this particular case.
The zeros of $F_{n}^{\mathsf{E}_{m}}$ for general $m$-gons are illustrated in [24] and appear to behave very similar to the ones for $T_{n}^{\mathsf{E}_{m}}$ when $m\geq 4$. We therefore believe that the zeros should approach the boundary when $m\geq 4$ in the sense that (11) should hold for every compact set in the interior. This would of course also imply that
@@ 4.3.1. Regular polygons
4.3.2. Circular lunes
Recall the definition of $\mathsf{C}_{\alpha}$ from (23). Based on the plot in Figure 6(a) it appears that most of the zeros approach the boundary in the case when $\alpha=1/2$ and that (11) should hold for any compact set contained in the interior. This is in fact a known result and follows from [44, Theorem 2.1]. Indeed, from there we gather that $\nu(T_{n}^{\mathsf{C}_{\alpha}})\xrightarrow{\ast}\mu_{\mathsf{C}_{\alpha}}$ as $n\rightarrow\infty$ for any $\alpha\in(0,1)$. In this sense, the computed polynomials serve to confirm the predicted behavior from theoretical results. For any value of $\alpha\in(0,1)\cup(1,2)$ it follows from [32, Theorem 1.5] that $\nu(F_{n}^{\mathsf{C}_{\alpha}})\xrightarrow{\ast}\mu_{\mathsf{C}_{\alpha}}$ along some subsequence. Motivated by the resemblance between the plots of zeros for Faber polynomials in [24] with the corresponding zeros of $T_{n}^{\mathsf{C}_{\alpha}}$ computed here, we suspect that